UV AND 532 NM PHOTODISSOCIATION OF O-NITROTOLUENE: DETECTION OF ELECTRONICALLY EXCITED NITRIC OXIDE IN NITROGEN AND ARGON

\underline{HELENA DIEZ-Y-RIEGA} and HERGEN EILERS, Applied Sciences Laboratory, Institute for Shock Physics, Washington State University, Spokane, WA 99210-1495, USA.

It is well known that NO is one of the main fragments produced by photodissociation of o-nitrotoluene. We detected vibrationally-excited NO in ground and electronically excited states using LIF. We also observed emission due to the formation of C(I) which is overlapping with NO emission. In the presence of N_{2}, longer lifetime of the NO emission is observed showing evidence of energy transfer from highly excited N_{2}. In the presence of Ar, evidence of o-nitrotoluene-Ar cluster formation was observed.